If one faucet is turned on alone the bucket will be filled in five minutes. Write the fraction of the bucket that will be filled in one minute?
Answer by Theo(3458)

r * t = u

r rate.
t = time
u = units

the unit is 1 bucket full of water.

the time is 5 minutes

the formula becomes:

r * 5 = 1

divide both sides of this equation by 5 to get:

r = 1/5

the bucket is being filled at the rate of 1/5 of the bucket per minute.

r * 1 = 1/5 * 1 = 1/5

in one minute, 1/5 of the bucket has been filled.